# Adaptive Batch Size for Privately Finding Second-Order Stationary Points

### v1

@@ 3.1 Constructing Private Gradient Oracles
By Proposition 3.4, setting $T=\tilde{O}(B/\eta\gamma^{2})$ suffices to find an $\tilde{O}(\gamma)$-SOSP. Let $\{x_{1},\cdots,x_{t}\}$ be the outputs of the algorithms, where $t\leq T$ denotes the step we halt the algorithm. We show
$\sum_{i=1}^{t}\|\widetilde{\nabla}_{i}\|_{2}\lesssim\tilde{O}(BM/\gamma).$ (2)
Denote the set $S:=\{i\in[t]:\|\widetilde{\nabla}_{i}\|\leq\gamma\}$. As $|S|\leq T=\tilde{O}(B/\eta\gamma^{2})$, we know $\sum_{i\in S}\|\widetilde{\nabla}_{i}\|\leq\tilde{O}(B/\eta\gamma)$.
Now consider the set $S^{c}:=[t]\setminus S$ denoting the index of steps when the norm of the gradient estimator is large. It suffices to bound $\sum_{i\in S^{c}}\|\widetilde{\nabla}_{i}\|_{2}$.
By Lemma 3.2, we know when ~$\|\widetilde{\nabla}_{i}\|\geq\gamma$, $F(x_{i+1})-F(x_{i})\leq-\eta\|\widetilde{\nabla}_{i}\|^{2}/16$, and when ~$\|\widetilde{\nabla}_{i}\|\leq\gamma,F(x_{i+1})\leq F(x_{i})+\eta\gamma^{2}$. Given the bound on the function values, we know
$\sum_{i\in S^{c}}\|\widetilde{\nabla}_{i}\|^{2}\leq\tilde{O}(B/\eta).$
@@
$\sum_{i\in S^{c}}\|\widetilde{\nabla}_{i}\|\leq\frac{\sum_{i\in S^{c}}\|\widetilde{\nabla}_{i}\|^{2}}{\gamma}\leq\tilde{O}(\frac{B}{\eta\gamma}).$
This completes the proof of Equation (2).
The total number of functions we used for $\mathcal{O}_{1}$ is upper bounded by $b\cdot\frac{\sum_{i\in[t]\|\widetilde{\nabla}_{i}\|}}{\kappa}=\tilde{O}(bBM/\kappa\gamma)$. The total number of functions we used for $\mathcal{O}_{2}$ is upper bounded as follows:
$\sum_{i\in[t]}b_{t}\lesssim(\frac{\sqrt{d}}{\alpha\varepsilon}+\frac{\kappa}{\alpha^{2}})\sum_{i\in[t]}\|\widetilde{\nabla}_{t}\|+T\leq BM\cdot\tilde{O}(\frac{\sqrt{d}}{\gamma^{2}\varepsilon}+\frac{\kappa}{\gamma^{3}}+\frac{1}{\gamma^{2}}).$
@@
Lemma 3.10.
$\alpha=O\Big{(}((BGM)^{1/3}+\sqrt{BM})(\frac{\sqrt{d}}{n\varepsilon})^{1/2}+\frac{B^{\frac{2}{9}}M^{\frac{2}{9}}G^{\frac{5}{9}}+B^{\frac{4}{9}}M^{\frac{4}{9}}G^{\frac{1}{9}}}{n^{1/3}}+\frac{(MB)^{1/2}}{\sqrt{n}}\Big{)},$
@@
$n\geq\tilde{\Omega}\big{(}\frac{bBM}{\kappa\gamma}+BM(\frac{\sqrt{d}}{\gamma^{2}\varepsilon}+\frac{\kappa}{\gamma^{3}}+\frac{1}{\gamma^{2}})\big{)}.$
@@
$n\geq\tilde{\Omega}(\frac{bBM}{\kappa\gamma})=\Theta(\frac{BGM\sqrt{d}}{\kappa\varepsilon\gamma^{2}}+\frac{G^{2}BM}{\kappa\gamma^{3}})\Leftarrow\gamma\geq\tilde{O}(\frac{(BGM)^{1/3}d^{1/4}}{\sqrt{n\varepsilon}}+\frac{B^{\frac{2}{9}}M^{\frac{2}{9}}G^{\frac{5}{9}}}{n^{1/3}}).$
@@
Theorem 3.11.
Given $\varepsilon,\delta>0$, with gradient oracles in Algorithm 3, setting $b=G(\sqrt{d}/\varepsilon\alpha+1/\alpha^{2}),b_{t}=\max\{\frac{\|\widetilde{\nabla}_{t}\|\cdot\sqrt{d}}{\alpha\varepsilon},\frac{\|\widetilde{\nabla}_{t}\|\kappa}{\alpha^{2}},1\}$, $\kappa=\max\{\frac{\alpha\sqrt{d}}{\varepsilon},(BGM)^{1/3}\}$ and $\sigma=\alpha/\sqrt{d}$, Algorithm 2 is $(\varepsilon,\delta)$-DP, and if the dataset is i.i.d. drawn from the underlying distribution $\mathcal{P}$, at least one of the output of is $\tilde{O}(\alpha)$-SOSP, where
$\alpha=O\Big{(}((BGM)^{1/3}+\sqrt{BM})(\frac{\sqrt{d}}{n\varepsilon})^{1/2}+\frac{B^{\frac{2}{9}}M^{\frac{2}{9}}G^{\frac{5}{9}}+B^{\frac{4}{9}}M^{\frac{4}{9}}G^{\frac{1}{9}}}{n^{1/3}}+\frac{(MB)^{1/2}}{\sqrt{n}}\Big{)}$
Remark 3.12.
If we treat the parameters $B,G,M$ as constants $O(1)$, then we get $\alpha=\tilde{O}((\frac{\sqrt{d}}{n\varepsilon})^{1/2}+\frac{1}{n^{1/3}})$ as claimed before in the abstract and introduction.
If we make further assumptions, like assuming the functions are defined over a constraint domain $\mathcal{X}\subset\mathbb{R}^{d}$ of diameter $R$ and we allow exponential running time, we can get some other standard bounds that can be better than Theorem 3.11 in some regimes. See Appendix A.2 for more discussions.
4 Discussion
Appendix A Appendix
A.1 Discussion of Proposition 3.4
Proposition 3.4 has become a standard result in non-convex optimization. Ganesh et al. [2023] assume two kinds of gradient oracles that satisfy the norm-Subgaussian definition below:
Definition A.1 (SubGaussian gradient oracles).
For a $G$-Lipschitz and $M$-smooth function $F$:
$(1)$ We say $\mathcal{O}_{1}$ is a first kind of $\zeta_{1}$ norm-subGaussian Gradient oracle if given $x\in\mathbb{R}^{d}$, $\mathcal{O}(x)$ satisfies $\operatorname*{\mathbb{E}}\mathcal{O}_{1}(x)=\nabla F(x)$ and $\mathcal{O}_{1}(x)-\nabla F(x)$ is $\mathrm{nSG}(\zeta_{1})$.
$(2)$ We say $\mathcal{O}_{2}$ is a second kind of $\zeta_{2}$ norm-subGaussian stochastic Gradient oracle if given $x,y\in\mathbb{R}^{d}$, $\mathcal{O}_{2}(x,y)$ satisfies that $\operatorname*{\mathbb{E}}\mathcal{O}_{2}(x,y)=\nabla F(x)-\nabla F(y)$ and $\mathcal{O}_{2}(x,y)-(\nabla F(x)-\nabla F(y))$ is $\mathrm{nSG}(\zeta_{2}\|x-y\|)$.
In Definition A.1, the oracles are required to be unbiased with norm-subGaussian error. While this condition is sufficient, it is not necessary. These two types of gradient oracles are then used to construct gradient estimators whose errors are tightly controlled with high probability (Lemma3.3 in Ganesh et al. [2023]). Proposition 3.4 can be established directly using Lemma 3.3 from Ganesh et al. [2023], as demonstrated in the proof of Lemma 3.6 in the same work. We include the proof here for completeness.
Proof of Proposition 3.4.
By Lemma 3.2 and the precondition that ~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\gamma$, we know that, if $\|\nabla F(x_{t})\|\geq 4\gamma$, then $F(x_{t+1})-F(x_{t})\leq-\eta\|\widetilde{\nabla}\|^{2}/16$. Otherwise, $\|\nabla F(x_{t})\|<4\gamma$. If $\|\nabla F(x_{t})\|<4\gamma$ but $x_{t}$ is a saddle point, then by Lemma 3.3, we know
where $\Gamma=\tilde{O}(M/\sqrt{\rho\gamma})$. Then if none of point in $\{x_{i}\}_{i\in[T]}$ is an $\tilde{O}(\gamma)$-SOSP, then we know $F(x_{T})-F(x_{0})<-B$, which is contradictory to Assumption 3.1. Hence at least one point in $\{x_{i}\}_{i\in[T]}$ should be an $\tilde{O}(\gamma)$-SOSP, and hence complete the proof.
A.2 Other results
@@
Theorem A.2.
Suppose $\mathcal{D}$ is i.i.d. drawn from the underlying distribution $\mathcal{P}$ and under Assumption 3.1. Additionaly assume $f(;\zeta):\mathcal{X}\to\mathbb{R}$ for some constrained domain $\mathcal{X}\subset\mathbb{R}^{d}$ of diameter $R$. Then we know for any point $x\in\mathcal{X}$, with probability at least $1-\iota$, we have
@@
The current SOTA of finding SOSP privately of $F_{\mathcal{D}}$ is from Ganesh et al. [2023], where they can find an $\tilde{O}((\sqrt{d}/n\varepsilon)^{2/3})$-SOSP. Combining the SOTA and Theorem A.2, we can find the $\alpha$-SOSP of $F_{\mathcal{P}}$ privately with

### v2

@@ 4 Discussion
By Proposition 3.10, setting $T=\tilde{O}(B/\eta\gamma^{2})$ suffices to find an $\tilde{O}(\gamma)$-SOSP. Let $\{x_{1},\cdots,x_{t}\}$ be the outputs of the algorithms, where $t\leq T$ denotes the step we halt the algorithm. We first show
$\sum_{i=1}^{t}\|g_{i}\|_{2}\lesssim\tilde{O}(BM/\gamma).$ (9)
Denote the set $S:=\{i\in[t]:\|g_{i}\|\leq\gamma\}$. As $|S|\leq T=\tilde{O}(B/\eta\gamma^{2})$, we know $\sum_{i\in S}\|g_{i}\|\leq\tilde{O}(B/\eta\gamma)$.
Now consider the set $S^{c}:=[t]\setminus S$ denoting the index of steps when the norm of the gradient estimator is large. It suffices to bound $\sum_{i\in S^{c}}\|g_{i}\|_{2}$.
By Lemma 3.2, we know when $\|g_{i}\|\geq\gamma$, $F(x_{i+1})-F(x_{i})\leq-\eta\|g_{i}\|^{2}/16$, and when $\|g_{i}\|\leq\gamma,F(x_{i+1})\leq F(x_{i})+\eta\gamma^{2}$. Given the bound on the function values, we know
$\sum_{i\in S^{c}}\|g_{i}\|^{2}\leq\tilde{O}(B/\eta).$
@@
$\sum_{i\in S^{c}}\|g_{i}\|\leq\frac{\sum_{i\in S^{c}}\|g_{i}\|^{2}}{\gamma}\leq\tilde{O}(\frac{B}{\eta\gamma}).$
This completes the proof of Equation (9). Moreover, we know the total time that we will escape from the saddle point is at most $O(B/\Phi)$. Hence, The total number of functions we used for $\mathcal{O}_{1}$ and $\mathcal{O}_{3}$, is upper bounded by
$b\cdot(\frac{\sum_{i\in[t]}\|g_{i}\|}{\kappa}+B/\Phi\tau)=\tilde{O}(bBM/\kappa\gamma).$
The total number of functions we used for $\mathcal{O}_{2}$ is upper bounded as follows:
$\sum_{i\in[t]}b_{t}\lesssim(\frac{\sqrt{d}}{\alpha\varepsilon}+\frac{\kappa}{\alpha^{2}}+\frac{\rho\kappa}{M^{2}\alpha})\sum_{i\in[t]}\|g_{i}\|+T\leq BM\cdot\tilde{O}(\frac{\sqrt{d}}{\gamma^{2}\varepsilon}+\frac{\kappa}{\gamma^{3}}+\frac{\rho\kappa}{M^{2}\gamma^{2}}+\frac{1}{\gamma^{2}}).$
@@
B.6 Proof of Lemma 3.16
@@
$n\geq\tilde{\Omega}\big{(}\frac{bBM}{\kappa\gamma}+BM(\frac{\sqrt{d}}{\gamma^{2}\varepsilon}+\frac{\kappa}{\gamma^{3}}+\frac{\rho\kappa}{M^{2}\gamma^{2}}+\frac{1}{\gamma^{2}})\big{)}.$
@@
$n\geq\tilde{\Omega}(\frac{bBM}{\kappa\gamma})=\Theta(\frac{BGM\sqrt{d}}{\kappa\varepsilon\gamma^{2}}+\frac{G^{2}BM}{\kappa\gamma^{3}}+\frac{BM^{2}\sqrt{d}}{\kappa\rho\gamma\varepsilon}+\frac{BM^{3}}{\kappa\rho\gamma^{2}})$
$\Leftarrow\gamma\geq\tilde{O}(\frac{(BGM)^{1/3}d^{1/4}}{\sqrt{n\varepsilon}}+\frac{B^{\frac{2}{9}}M^{\frac{2}{9}}G^{\frac{5}{9}}}{n^{1/3}}+\frac{\sqrt{BM^{2}}}{\sqrt{\rho n}}+\frac{B^{1/3}M^{4/3}}{G^{1/6}\sqrt{\rho n}}).$
@@
Fourthly,
$n\geq\tilde{\Omega}(\frac{B\rho\kappa}{M\gamma^{2}})\geq\tilde{\Omega}(\frac{B^{4/3}G^{1/3}\rho}{M^{2/3}\gamma^{2}}+\frac{B\rho\sqrt{d}}{M\gamma\varepsilon})$
$\Leftarrow\gamma\geq\tilde{O}(\frac{B^{2/3}G^{1/6}\sqrt{\rho}}{M^{1/3}\sqrt{n}}+\frac{B\rho\sqrt{d}}{Mn\varepsilon}).$
@@
Appendix C Other results
@@
Theorem C.1.
Suppose $\mathcal{D}$ is i.i.d. drawn from the underlying distribution $\mathcal{P}$ and under Assumption 3.1. Additionally assume $f(;\zeta):\mathcal{X}\to\mathbb{R}$ for some constrained domain $\mathcal{X}\subset\mathbb{R}^{d}$ of diameter $R$. Then we know for any point $x\in\mathcal{X}$, with probability at least $1-\iota$, we have
@@
The current SOTA of finding SOSP privately of $F_{\mathcal{D}}$ is from Ganesh et al. [2023], where they can find an $\tilde{O}((\sqrt{d}/n\varepsilon)^{2/3})$-SOSP. Combining the SOTA and Theorem C.1, we can find the $\alpha$-SOSP of $F_{\mathcal{P}}$ privately with